OpenStudy (anonymous):

can you show me how to factor this!!!!! will give medal!! need help with 2

OpenStudy (anonymous):

2x^2 + 13x + 15 2x^2 + 27x + 13

OpenStudy (anonymous):

2x^2 + 10x + 3x + 15 (2x+3)(x+5) im thinking this is wat u do for the first one

OpenStudy (anonymous):

First one in factored form is: (5 + x) (3 + 2 x) Second one in factored form is: (13 + x) (1 + 2 x) Hope I helped!

OpenStudy (anonymous):

@4kec4 can you expalin to me how to do the first and second?

OpenStudy (anonymous):

ok first i set it up like this:\[\left( 2x+... \right)\left( x+... \right)\] then figure out what two numbers give a product of 15. In this case 3*5=15. when you put 3 and 5 in you see that:\[\left( 2x+3 \right)\left( x+5 \right)=2x^2+10x+3x+15=2x^2+13x+15\] Remember you multiply like this: